OpenStudy (anonymous):

how to find the exact value of cos 7pi/4

OpenStudy (anonymous):

*without a calculator

OpenStudy (anonymous):

pi=180 degrees.

OpenStudy (anonymous):

7*180=?

OpenStudy (whpalmer4):

2 * pi radians is 1 whole trip around the unit circle. cos 7pi/4 = - cos pi/4. Can you draw the triangle in the unit circle for an angle of pi/4? you should be able to work out the rest with the Pythagorean theorem.

OpenStudy (anonymous):

cos(7pi/4) =cos[(8-1)pi/4] =cos[8pi/4-1pi/4] =cos[2pi-pi/4] =cos(pi/4) =cos(180/4) =cos(45) now you say the value.